# with theoretical and observational applications

### v1

@@ 4.4.3 $w_{0}w_{a}$ parametrisation and phantom behaviour
One analytical parametrisation of $w_{\varphi}$ in the matter - dark energy phase is the Chevallier-Polarski-Linder (CPL) or $w_{0}w_{a}$ parametrisation. It proposes a linear behaviour in $a$ around today ($a=1$)
@@
In quintessence models and with our analytical solutions, one starts around $N_{m}$ with $w_{\varphi}\approx-1$. With this as an anchor point, if the line aims at a small $|w_{0}|$, then the slope $|w_{a}|$ would be large; on the contrary, we can also have a large $|w_{0}|$ and a small $|w_{a}|$. Figure 23 illustrates these two possibilities.
The DESI values are difficult to match in general, because the central values of $|w_{0}|$ and $|w_{a}|$ are both large. [^5] Planck 2018 results [26] discuss as well possible values for $w_{0},w_{a}$ using several data sets. Two data sets are presented in [26, Tab. 6] with equally good fits. One is highlighted in the main text, because it gives more precise results and can be compatible with no phantom behaviour: it provides central values close $\Lambda$CDM, namely $w_{0}=-0.957,w_{a}=-0.29$. Allowing however some phantom behaviour, we can also consider the other data set results, giving $w_{0}=-0.76,w_{a}=-0.72$. Interestingly, the latter are close to the DESI values. Such values (and parameters) imply that $w_{\varphi}<-1$ at $N_{m}$, contrary to the above. Values such that $w_{\varphi}<-1$ are referred to as a phantom behaviour. As discussed in Section 2.1, a quintessence model admits by definition $-1\leq w_{\varphi}\leq 1$ so a priori, it cannot lead to a phantom behaviour, hence the difficulties encountered in reproducing DESI results. We illustrate this situation in Figure 24 comparing DESI results to our hilltop example.
@@
Let us point-out that a quintessence model can lead to some phantom behaviour as an artefact of the $w_{0}w_{a}$ parametrisation. Indeed, as mentioned, one has $w_{\varphi}\approx-1$ at matter domination and for some time afterwards, and then $w_{\varphi}$ starts growing. Therefore, $\partial_{a}w_{\varphi}$ was initially vanishing and then becomes positive, from which we deduce that $\partial_{a}^{2}w_{\varphi}>0$. This shows that the initial growth is convex. [^6] Note that for a function $f$, $\partial_{N}f=e^{N}\,\partial_{a}f$, meaning that growing in $a$ is equivalent to growing in $N$. However, one has $\partial_{N}^{2}f=e^{2N}\,\partial_{a}^{2}f+e^{N}\,\partial_{a}f$. This means that convex in $N$, as observed in our two examples, does not imply convex in $a$. A counter-example to such an equivalence is indeed provided in Figure 23(a) by the exponential quintessence. The best fit line of a convex curve necessarily crosses it. Therefore, this fitting line must initially be below the curve, at least if the latter stays convex. For a curve which is at $-1$ for some time before rising, then the best fit line must go below $-1$: this generates a (fake) phantom behaviour. This parametrisation artefact is visible in Figure 23(b), where the red line goes below $-1$ at the first moments considered; it can be made even more manifest by increasing the potential parameter $k$ in that model.
@@
Let us add that a DESI “model-agnostic” reconstruction of $w_{\varphi}(z)$ [36, Fig. 1] shows again large portion of phantom behaviour in the past (although DESI + Planck data alone is not displayed); equivalently, $\rho_{\varphi}/\rho_{\varphi 0}$ (corresponding in that work to $f_{DE}$) is shown to grow with time towards the past. The latter is not possible within quintessence because of Hubble friction, as indicated by (2.6). With (2.7), this is equivalent to say that one must have $w_{\varphi}\geq-1$. Insisting on having such a phantom behaviour, that goes beyond an artefact of the $w_{0}w_{a}$ parametrisation, would thus be a major theoretical challenge, requiring for example non-standard (negative) kinetic energy, non-minimal couplings or modified gravity (see e.g. [26, Sec. 7.4.1]).
A possible way-out can be seen in the fact that error bars quickly become huge for $w_{\varphi}(z)$ in [36] once it becomes phantom. This is related to the fact that the measurement is challenging in the past, especially when it is not dark energy dominated. Therefore one option would be to discard the data (and period of time) indicating a phantom behaviour, and wait for more precise measurements in that earlier phase. Doing so may allow a match to a quintessence model with the most recent period of time.
@@ 4.4.4 Bounds on $w_{0},w_{a}$
$\text{{\bf Lemma.}}\quad\quad\boxed{\hskip 10.00002pt\exists\ a>0\ {\rm s.t.}\ w_{\varphi}<-1\quad\Leftrightarrow\quad w_{0}+w_{a}<-1\hskip 10.00002pt}$ (4.67)
assuming the $w_{0}w_{a}$ parametrisation to hold as well as a growing $w_{\varphi}$ (meaning $w_{a}<0$).
@@ 5 Outlook: string theory, cosmology, and observational targets
Another comment has to do the electroweak phase transition in the early universe, during which the Higgs field is supposed to roll-down its potential (while the potential evolves thermally to generate a new minimum). This “rolling-down” classical picture is probably only valid for a second order phase transition; a first order one may rather happen by a non-perturbative creation of bubbles. In any case, this phase transition may be considered to happen during radiation domination: if that is the case, the freezing described above could affect the Higgs scalar field, at least for a small enough scalar potential. This would prevent the naive rolling-down and thus possibly forbid a (second order) phase transition.
@@
    The first obvious target, already tackled by current and coming experiments, is the recent evolution of $w_{\varphi}$, away from $-1$. As stressed in our study, the quintessence models considered generically give $w_{\varphi}\approx-1$ at matter domination ($N\approx 2.5$, $z\approx 11$), but also before that, and after that until $N=N_{c}$. We suggested that $N_{c}\geq-1.5$, giving $z_{c}\leq 3.5$, depending on the model. The fact that $w_{\varphi}\approx-1$ for so long makes it difficult to distinguish the resulting evolution from that of $\Lambda$CDM. We combined this feature together with the $w_{0}w_{a}$ parametrisation into a schematic parametrisation described in Figure 25, that could avoid the phantom behaviour issues.
@@ Acknowledgements
We would like to thank G. Bélanger, J. Larena, A. Moradinezhad, V. Poulin and P. Salati for insightful discussions.

### v2

@@ 4.4.3 $w_{0}w_{a}$ parametrisation and phantom behaviour
One analytic parametrisation of $w_{\varphi}$ in the matter - dark energy phase is the Chevallier-Polarski-Linder (CPL) or $w_{0}w_{a}$ parametrisation. It proposes a linear behaviour in $a$ around today ($a=1$)
@@
In quintessence models and with our analytic solutions, one starts around $N_{m}$ with $w_{\varphi}\approx-1$. With this as an anchor point, if the line aims at a small $|w_{0}|$, then the slope $|w_{a}|$ would be large; on the contrary, we can also have a large $|w_{0}|$ and a small $|w_{a}|$. Figure 23 illustrates these two possibilities.
The DESI values are difficult to match in general, because the central values of $|w_{0}|$ and $|w_{a}|$ are both large. [^5] Planck 2018 results [30] discuss as well possible values for $w_{0},w_{a}$ using several data sets. Two data sets are presented in [30, Tab. 6] with equally good fits. One is highlighted in the main text, because it gives more precise results and can be compatible with no phantom behaviour: it provides central values close $\Lambda$CDM, namely $w_{0}=-0.957,w_{a}=-0.29$. Allowing however some phantom behaviour, we can also consider the other data set results, giving $w_{0}=-0.76,w_{a}=-0.72$. Interestingly, the latter are close to the DESI values. Such values (and parameters) imply that $w_{\varphi}<-1$ at $N_{m}$, contrary to the above. Values such that $w_{\varphi}<-1$ are referred to as a phantom behaviour. As discussed in Section 2.1, a quintessence model admits by definition $-1\leq w_{\varphi}\leq 1$ so a priori, it cannot lead to a phantom behaviour, hence the difficulties encountered in reproducing DESI results. We illustrate this situation in Figure 24 comparing DESI results to our hilltop example.
@@
Let us point-out that a quintessence model can lead to some phantom behaviour as an artefact of the $w_{0}w_{a}$ parametrisation (see [42] for related ideas with the hilltop potential). Indeed, as mentioned, one has $w_{\varphi}\approx-1$ at matter domination and for some time afterwards, and then $w_{\varphi}$ starts growing. Therefore, $\partial_{a}w_{\varphi}$ was initially vanishing and then becomes positive, from which we deduce that $\partial_{a}^{2}w_{\varphi}>0$. This shows that the initial growth is convex. [^6] Note that for a function $f$, $\partial_{N}f=e^{N}\,\partial_{a}f$, meaning that growing in $a$ is equivalent to growing in $N$. However, one has $\partial_{N}^{2}f=e^{2N}\,\partial_{a}^{2}f+e^{N}\,\partial_{a}f$. This means that convex in $N$, as observed in our two examples, does not imply convex in $a$. A counter-example to such an equivalence is indeed provided in Figure 23(a) by the exponential quintessence. The best fit line of a convex curve necessarily crosses it. Therefore, this fitting line must initially be below the curve, at least if the latter stays convex. For a curve which is at $-1$ for some time before rising, then the best fit line must go below $-1$: this generates a (fake) phantom behaviour. This parametrisation artefact is visible in Figure 23(b), where the red line goes below $-1$ at the first moments considered; it can be made even more manifest by increasing the potential parameter $k$ in that model.
@@
Let us add that a DESI “model-agnostic” reconstruction of $w_{\varphi}(z)$ [43, Fig. 1] shows again large portion of phantom behaviour in the past (although DESI + Planck data alone is not displayed); equivalently, $\rho_{\varphi}/\rho_{\varphi 0}$ (corresponding in that work to $f_{DE}$) is shown to grow with time towards the past. The latter is not possible within quintessence because of Hubble friction, as indicated by (2.6). With (2.7), this is equivalent to say that one must have $w_{\varphi}\geq-1$. Insisting on having such a phantom behaviour, that goes beyond an artefact of the $w_{0}w_{a}$ parametrisation, would thus be a major theoretical challenge, requiring for example non-standard (negative) kinetic energy, non-minimal couplings or modified gravity (see e.g. [30, Sec. 7.4.1]).
A possible way-out can be seen in the fact that error bars quickly become huge for $w_{\varphi}(z)$ in [43] once it becomes phantom. This is related to the fact that the measurement is challenging in the past, especially when it is not dark energy dominated. Therefore one option would be to discard the data (and period of time) indicating a phantom behaviour, and wait for more precise measurements in that earlier phase. Doing so may allow a match to a quintessence model with the most recent period of time. Further interesting remarks on the reality of such a phantom regime and the observations can be found in [44, App. A].
@@ 4.4.4 Bounds on $w_{0},w_{a}$
[…]
assuming the $w_{0}w_{a}$ parametrisation to hold as well as a growing $w_{\varphi}$ (meaning $w_{a}<0$). For completeness, let us mention that a related implication (not equivalence) was discussed in [45].
@@ 5 Outlook: string theory, cosmology, and observational targets
Another comment has to do the electroweak phase transition in the early universe, during which the Higgs field is supposed to roll-down its potential (while the potential evolves thermally to generate a new minimum) [54, 55]. This “rolling-down” classical picture is probably only valid for a second order phase transition; a first order one may rather happen by a non-perturbative creation of bubbles. In any case, this phase transition may be considered to happen during radiation domination: if that is the case, the freezing described above could affect the Higgs scalar field, at least for a small enough scalar potential. This would prevent the naive rolling-down and thus possibly forbid a (second order) phase transition.
@@
    The first obvious target, already tackled by current and coming experiments, is the recent evolution of $w_{\varphi}$, away from $-1$. As stressed in our study, the quintessence models considered generically give $w_{\varphi}\approx-1$ at matter domination ($N\approx-2.5$, $z\approx 11$), but also before that, and after that until $N=N_{c}$. We suggested that $N_{c}\geq-1.5$, giving $z_{c}\leq 3.5$, depending on the model. The fact that $w_{\varphi}\approx-1$ for so long makes it difficult to distinguish the resulting evolution from that of $\Lambda$CDM. We combined this feature together with the $w_{0}w_{a}$ parametrisation into a schematic parametrisation described in Figure 25, that could avoid the phantom behaviour issues.
@@ Acknowledgements
We would like to thank G. Bélanger, J. Larena, A. Moradinezhad, V. Poulin, F. Revello, P. Salati and F. Tonioni for insightful discussions.